OpenStudy (anonymous):

∠PQR and ∠ABC are supplementary angles and ∠PQR is eleven times as large as ∠ABC. Determine the measure of each angle.

OpenStudy (anonymous):

x+y =180 x=11y

OpenStudy (anonymous):

x= angle PQR y= angle ABC

OpenStudy (anonymous):

solve simultaneously

OpenStudy (anonymous):

no

OpenStudy (anonymous):

no 11x+y=180

OpenStudy (anonymous):

still wrong

OpenStudy (anonymous):

sub the second into the first 11y +y =180 , 12y = 180

OpenStudy (anonymous):

15

OpenStudy (anonymous):

y=15 sub into eqn 2 , x =11 x 15 = 165

OpenStudy (anonymous):

which one is the 2nd eqn?